can someone simplify (4b)^-3 for me? and show me how you did it?
Answer by stanbon(75887)

simplify (4b)^-3
----
The negative means "invert"
The 3 means "cube"
-----
(4b)^-3 = (1/(4b))^3 = 1/(64b^3)
================================
Cheers,
Stan H.
================================